Since interest rates have dropped, you consider refinancing your mortgage at a lower 6% rate.

If you took out a new 30-year mortgage at 6% for your remaining loan balance, what would your new monthly payments be?
1
Expert's answer
2019-12-10T11:00:52-0500

Let The Loan Amount = $160,000

Let the down payment = 10%

Given,

Interest rate = 6% per year = "\frac {6}{12} \space per \space month= 0.5 \% per \space month = \frac {0.5}{100} = 0.005"


Number of months = "30 \times 12 = 360"

Down payment = "10\% \space of \space 160000 = \$16000"


Loan amount = $160,000 - $16000 = $144000

Monthly payments will be calculated using the following formula


Monthly payment =

"\frac {P \times R \times (1 +R)^n} {(1+R)^{n-1}}"

Plug the values in this formula, then we have


Monthly payment =

"\frac {144000 \times 0.005 \times (1 +0.005)^{360}} {(1+0.005)^{360}-1}"



"=\$863"

Answer:


New monthly payments be $863